Consider gallium arsenide at T=300 K, assume the photon intensity at a particular point is I(x)=

0.05 W/cm2 at a wavelength of 0.75 m. The absorption coefficient at this wavelength is 0.9x104

cm-1

.

(a) Calculate the generation rate of electron hole pairs.

(b) Calculate the steady state excess concentration of electrons if the minority carrier (hole)

life time is 10-7

sec.


1
Expert's answer
2021-12-24T11:31:38-0500

the generation rate is given by:

"T = 300K"

"P_l=5*10^{-4}\large\frac{W}{m^2}"

"\lambda = 0.75m"

"\alpha = 0.9*10^6 m^{-1}"

"G -?"

"G = Q_e*\large\frac{\alpha P_l}{ħω} = \alpha P_l*e^{\alpha x} = 9 * 10^5*5*10^{-4}*e^{-0.75*9*10^5}"